Kona arrived at ISB at 9 am. The collector doors were closed and the fog pi system had crashed; no fog water collected. She rinsed the collecting bottle and frames with DI water at the sink, spray rinsed the rest of the apparatus, ran a test run and activated the system. Ready for fog collection.
New RH Sensor installed with Dan Fernandez as well as installation of new Raspberry Pi. Got a full jar sample that morning. hoping to have WiFi connection by monday or early next week
The collector did not open last night, and the RH at the weather station only hit about 70%. So I think we are good. No other actions taken, just waiting for fog now.
Today Celeste D. attempted the first deep clean at Pepperwood, but could not fit strand frames into 5g bucket. So proceeded to clean them with 409 as per SOP. This was the first time the collector was deployed since replacing the power source, fan switch, relay and wires (it has been dry at Pepperwood!). The RH was reading 33% at 4pm and matched the adjacent weather stations readings. I also noticed a few juvenile cockroaches in the collector and sample bottle. They are probably getting in thorugh very small spaces between the plexi and the teflon tray or where the doors meet the collector.
Celeste arrived at 7:00am today to an empty bottle once again. 😐 Fog was reportedly not heavy last night, but curiously the collector was open, RH was at 93.3% and the big fan was not on. I stopped the collector, then reactivated it and it fired up fine. At 7:30am the RH dipped below 90% and the fan turned off.
There was A LOT of heat in the electrical box upon arrival and the heat was apparently radiating from the wire connecting Big Fan Relay to the Big Fan Switch (the same wire that burned out previously). It was hot! Maybe this relates to the fan not being on during RH>90?
Jeremiah arrived at the ISB Greenhouse at 8:45 am. There was no fog sample collected and the collector doors had opened. He removed the strand frames, teflon tray, teflon side pieces, top plate, and drip tube assembly. He washed them with 409 and water and gave a DI rinse. He reassembled and took a blank sample (B-UCSC-140624). The system was reset and ready to go for sampling tonight.
Kona and Peter met at the LML site at 10:30 am. No fog was collected, although the doors had opened. We removed the strand frames and washed with DI at the sink. We rinsed the collector body and drip tray/tube assembly in place with copious amounts of DI water. Reassembled and reset for sampling tonight.
Installed both active and passive systems today. Active collector set up and ready to collect sample. Soaked collector components in 10% acid bath on 6/22 overnight.
Cloudy overnight, overcast was not at ground level. No fog sample collected. System washed by removing strand frames and rinsing with DI water at tap. Used sprayer on Teflon tray still in place. Reset system.